A test shows that 10 out of 100 gadgets are defective. In a random sample of two gadgets are chosen, the probability that both are defective is?
a. 1/1000
b. 1/110
c. 1/11
d. 1/10

Answer by stanbon(75887)

A test shows that 10 out of 100 gadgets are defective. In a random sample of two gadgets are chosen, the probability that both are defective is?
-----------
# of ways to pick 2 out of the 10 = 10C2 = [10*9]/[1*2] = 45
# of ways to pick 2 out of the 100=100C2 = [100*99]/[1*2] = 4950
---------------
P(pick 2 defective) = 45/4950 = 0.0091 = 1/110
=============================
Cheers,
Stan H.